OpenStudy (watermelon14):

find the exact value of cot(9pi/8)

OpenStudy (watermelon14):

@jim_thompson5910 ?

OpenStudy (watermelon14):

i know this requires the half angle formula

jimthompson5910 (jim_thompson5910):

first I would try to find tan(9pi/8)

OpenStudy (watermelon14):

\[1-\cos \theta/\sin\]

jimthompson5910 (jim_thompson5910):

http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf look on page 2 where it says "Half Angle Formulas "

OpenStudy (watermelon14):

oh isn't my formula correct?

jimthompson5910 (jim_thompson5910):

I would use \[\Large \tan\left(\frac{\theta}{2}\right) = \pm\sqrt{\frac{1-\cos(\theta)}{1+\cos(\theta)}}\]

jimthompson5910 (jim_thompson5910):

theta in this case is theta = 9pi/4

jimthompson5910 (jim_thompson5910):

notice how 9pi/4 - 2pi = 9pi/4 - 8pi/4 = pi/4 so 9pi/4 and pi/4 are coterminal

jimthompson5910 (jim_thompson5910):

so you could use theta = pi/4

OpenStudy (watermelon14):

ok and when it is 9pi/8 wouldn't you flip it and mult it by 2?

OpenStudy (watermelon14):

then plug that in the tan formula

jimthompson5910 (jim_thompson5910):

focus on pi/4 for now. Plug that into the formula I gave you and tell me what you get

OpenStudy (watermelon14):

ok but when u subtracted by 2pi. where did 2pi come from

OpenStudy (watermelon14):

oh because coterminal?

jimthompson5910 (jim_thompson5910):

yes adding or subtracting 2pi to any angle leads to another coterminal angle

OpenStudy (watermelon14):

is it \[\sqrt{4-\pi}\]

jimthompson5910 (jim_thompson5910):

what is cos(pi/4) equal to?

OpenStudy (watermelon14):

\[\sqrt{2}/2\]

jimthompson5910 (jim_thompson5910):

So you'd have \[\Large \tan\left(\frac{\theta}{2}\right) = \pm\sqrt{\frac{1-\cos(\theta)}{1+\cos(\theta)}}\] \[\Large \tan\left(\frac{\frac{\pi}{4}}{2}\right) = \pm\sqrt{\frac{1-\cos\left(\frac{\pi}{4}\right)}{1+\cos\left(\frac{\pi}{4}\right)}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \pm\sqrt{\frac{1-\frac{\sqrt{2}}{2}}{1+\frac{\sqrt{2}}{2}}}\]

jimthompson5910 (jim_thompson5910):

now simplify

OpenStudy (watermelon14):

\[\sqrt{3-2\sqrt{2}}\]

jimthompson5910 (jim_thompson5910):

notice how using a calc, we get tan(pi/8) = 0.4142135623731 so we can drop the minus in the plus/minus to get just this \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{1-\frac{\sqrt{2}}{2}}{1+\frac{\sqrt{2}}{2}}}\]

jimthompson5910 (jim_thompson5910):

now let's simplify the right side

jimthompson5910 (jim_thompson5910):

\[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{1-\frac{\sqrt{2}}{2}}{1+\frac{\sqrt{2}}{2}}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{\frac{2}{2}-\frac{\sqrt{2}}{2}}{\frac{2}{2}+\frac{\sqrt{2}}{2}}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{\frac{2-\sqrt{2}}{2}}{\frac{2+\sqrt{2}}{2}}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{2-\sqrt{2}}{2} \times \frac{2}{2+\sqrt{2}}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{2-\sqrt{2}}{2+\sqrt{2}}}\] Hopefully this is making sense?

jimthompson5910 (jim_thompson5910):

Now let's rationalize the denominator \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{2-\sqrt{2}}{2+\sqrt{2}}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{2-\sqrt{2}}{2+\sqrt{2}}\times\frac{2-\sqrt{2}}{2-\sqrt{2}}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{(2-\sqrt{2})(2-\sqrt{2})}{(2+\sqrt{2})(2-\sqrt{2})}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{4-4\sqrt{2}+2}{(2)^2-(\sqrt{2})^2}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{4-4\sqrt{2}+2}{4-2}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{4-4\sqrt{2}+2}{2}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{\frac{6-4\sqrt{2}}{2}}\] \[\Large \tan\left(\frac{\pi}{8}\right) = \sqrt{3-2\sqrt{2}}\]